7.What is the distance between the points x and y, given that x = 3 + 2i and y = 4 + 3i?
a.2
b.square root of 2
c.square root of 3
d.3

8.If r is the hypothenus of a right-angled triangle with sides a and b and if i and j are unit vectors in the horizontal and vertical directions, then r is given by
a.a^2i+b^2j
b.i + j
c.ai–bj
d.ai + bj
1
Expert's answer
2019-04-11T10:36:50-0400

"7. \; d=\sqrt{(3-4)^2+(2-3)^2}=\sqrt{2}\;\; Var.b.\\
8.\; r=ai+bj. \;\; Var. d."